### v1

# QCRD: Quality-guided Contrastive Rationale Distillation for Large Language Models

Abstract
Deploying large language models (LLMs) poses challenges in terms of resource limitations and inference efficiency. To address these challenges, recent research has focused on using smaller task-specific language models, which are enhanced by distilling the knowledge rationales generated by LLMs. However, previous works mostly emphasize the effectiveness of positive knowledge, while overlooking the knowledge noise and the exploration of negative knowledge. In this paper, we first propose a general approach called quality-guided contrastive rationale distillation for reasoning capacity learning, considering contrastive learning perspectives. For the learning of positive knowledge, we collect positive rationales through self-consistency to denoise the LLM rationales generated by temperature sampling. For the negative knowledge distillation, we generate negative rationales using temperature sampling for the iteration-before smaller language models themselves. Finally, a contrastive loss is designed to better distill the positive and negative rationales into the smaller language model, where an online-update discriminator is used to judge the qualities of rationales and assign weights for better optimizing the training process. Through extensive experiments on multiple reasoning tasks, we demonstrate that our method consistently outperforms the previous distillation methods and produces higher-quality rationales.
1 Introduction
The reasoning capabilities of large language models (LLMs) have been observed to scale with their model sizes, while necessitating substantial memory and computing resources Chowdhery et al. (2023); Wei et al. (2022a). As such, efficient model compression is crucial in the deployment of LLMs, especially on resource-limited devices or platforms. In recent years, knowledge distillation from an LLM to a smaller language model has emerged as a powerful and promising technique for model compression Hinton et al. (2015); Phuong and Lampert (2019). However, it is still open how to best close the performance gap with respect to LLMs on complex reasoning tasks Zelikman et al. (2022).
In this regard, it has more recently been shown that adding explanation-augmented prompts, especially, Chain-of-Thought (CoT) Wei et al. (2022b), can enable LLMs to generate reasonable explanations to justify the reasoning outcomes Li et al. (2022). In addition, distilling these rationales into smaller language models has been demonstrated to effectively improve the overall performance Hsieh et al. (2023); Li et al. (2022). Specifically, distilling Step-by-Step (DSS) Hsieh et al. (2023) was proposed as an innovative CoT distillation approach, which employed rationales from an LLM to guide a smaller language model under a multi-task learning setting. The approach involved training the smaller language model simultaneously on both label prediction and rationale generation tasks, effectively leveraging their mutual benefits.
Figure 1: Comparison between previous methods and our proposed method, where circle points denote rationales, and colors of the circle points correspond to rationale types, and shades of darker indicate higher qualities. The "align" means minimizing the distance between rationales, while the "repel" means maximizing the distance.
The essence of distilling rationales is to guide the model in learning additional knowledge related to the labels, and the knowledge can be concluded into two class: positive and negative. Previous works, although effective, still suffer from some drawbacks. On the one hand, the positive knowledge for distillation is noisy. Methods Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022) treat the rationales generated by LLMs as golden answers and aim to minimize the gap between these rationales and those generated by smaller language models. However, despite LLM’s powerful zero-shot/few-shot abilities, it may occasionally produce incorrect reasoning steps, leading to erroneous rationales/answers. Such erroneous rationales may degrade the reasoning performance of the distilled smaller language models. On the other hand, generating negative rationales and incorporating them into CoT distillation remain understudied, while negative knowledge has early proved to be constructive and effective for models.
To this end, we propose a general method, named Quality guided Contrastive Rationale Distillation (QCRD), to guide the knowledge distillation to a smaller language model from a contrastive learning perspective. The comparison between the previous methods and our proposed QCRD is illustrated in Fig. 1. Specifically, the previous methods focus on the alignment between the rationale of the student model and the corresponding one of the teacher model, while our proposed QCRD aligns the student’s distribution and contrastive knowledge distribution with different sampled rationales. The core design of QCRD is to generate a diverse set of contrastive rationales and efficiently distill them into the student model. For the positive part, to ensure the qualities of positive rationales, we first prompt the LLM and sample the output to generate multi-round rationales for each input question. We then apply the self-consistence to denoise the rationale set and split it into positive samples and negative samples. For the negative part, we employ a self-play strategy Silver et al. (2018) during the training to generate low-quality rationales from previous iterations of the smaller language model and treat them as negative samples. Finally, for better knowledge learning, we present a contrastive loss to distill both the positive and negative rationales into the smaller language model. And a discriminator is adopted to assess the qualities of the rationales and assign them appropriate weights to optimize the training process across the datasets. To demonstrate the superiority of QCRD, we conduct comprehensive experiments with two smaller types of T5 models Raffel et al. (2020), i.e., T5-base (220M parameters) and T5-small (60M parameters), on four popular datasets, followed by detailed analysis and discussion.
Our main contributions of this paper can be summarized below.
  • We first develop a general CoT distillation approach (i.e., QCRD) from a contrastive learning perspective, aiming to guide the student model to learn both positive and negative knowledge from rationales.
  • We explore a contrastive distillation loss to facilitate effective distillation of the generated positive and negative rationales, where the qualities of the rationales judged by a discriminator are considered to optimize the training process across the whole datasets.
  • Experimental results across multiple datasets show that QCRD demonstrably outperforms existing benchmarks, evidencing its efficiency in transferring contrastive reasoning knowledge to the smaller language model.
Figure 2: Illustration of the proposed quality guided contrastive rationale distillation (QCRD) for distilling contrastive knowledge from teacher models into the student model, where CE denotes the cross entropy.
2 Related Work
Our work utilizes the emerging reasoning capabilities of current LLMs to distill task-specific knowledge into smaller specialist models. We present a brief overview of previous studies across four areas related to our work in the following.
Knowledge distillation from LLMs. Knowledge distillation (KD) is a highly effective technique for transferring knowledge from larger, more capable teacher models to smaller student models that are more suitable for practical applications Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022). The technique can be generally classified into two different categories, namely, Black-box KD and White-box KD. White-box KD provides access to the teacher’s parameters, enabling the student LM to gain a deeper understanding of the teacher LLM’s internal structure and knowledge representations, while only the teacher’s predictions are accessible in Black-box KD. Both of Black-box and White-box KDs have shown promising results in fine-tuning small models on the prompt response pairs generated by LLMs Zhu et al. (2023). In this paper, we hypothesize that only the prediction labels and rationales generated by LLMs are accessible.
Multi-task learning with LLM generated rationales. Current LLMs have already shown their capabilities to generate high-quality reasoning steps, resulting in explanations of their predictions Kojima et al. (2022), and these explanations have been found to be valuable additional knowledge for fine-tuning smaller models Hsieh et al. (2023). In order to empower smaller models with robust reasoning abilities and the capacity to generate explanations, a multi-task learning framework is commonly employed. This framework enables the integration of strong reasoning capabilities into smaller models while maintaining their original functionality. However, previous approaches have applied a single loss form, which only focuses on aligning the output of the small model with that of the LLM Hsieh et al. (2023); Magister et al. (2022).
Self-consistency of LLMs. The self-consistency of LLMs refers to the capacity to maintain coherent and rational reasoning during input processing. Based on the intuition that complex reasoning tasks typically admit multiple reasoning paths that reach a correct answer, the self-consistency can improve the LLMs’ reasoning performance by integratedly sampling CoT outputs and choosing the most consistent answer Stanovich and West (1991). The self-consistency has proven to be an effective technique for denoising datasets and selecting positive and negative samples Wang et al. (2022).
Contrastive learning for LLMs. Contrastive learning has demonstrated its efficiency across diverse domains, encompassing computer vision, natural language processing, and recommendation systems Jaiswal et al. (2020); Le-Khac et al. (2020). Notably, the application of contrastive learning to LLMs has emerged as a recent development, highlighting the effectiveness of incorporating negative samples and showing promising outcomes Li et al. (2024). However, to the best of our knowledge, the application of contrastive learning in CoT distillation has not been explored thus far.
3 Methodology
In this paper, we first propose a general contrastive CoT distillation approach, i.e., quality guided contrastive rationale distillation (QCRD), for training smaller models by distilling contrastive knowledge from teacher models. Our approach leverages the reasoning capabilities and self-consistency of LLMs, as well as the self-supervised ability of smaller models themselves. As illustrated in Fig. 2, our approach consists of the following three parts. (1) Following the method described in Hsieh et al. (2023), we apply a multi-task framework for the supervised training of the student model, where the label task aligns the prediction label with the ground truth and the rationale task learns the knowledge from the generated rationales, as can be seen in Fig. 2a. (2) As displayed in Fig. 2b, we design a general approach to generate contrastive knowledge. Specifically, for the positive knowledge, we prompt the LLM to generate multi-round output labels and corresponding rationales using temperature sampling. Collected rationale samples are denoised based on the ground truth or the majority of the self-consistency results. For the negative knowledge, we consider the remaining noised rationales from the LLM as negative samples. Moreover, we generate additional lower-quality rationales by iterating the smaller model itself. These lower-quality rationales are thought of as negative samples. (3) As shown in in Fig. 2c, for better knowledge learning from rationales, We design a quality-guided contrastive learning strategy, where a contrastive loss is applied with the guidance of an online-updated discriminator to distinguish between positive and negative samples and assign them quality scores.
3.1 Multi-task framework for the student model
Previous works have already demonstrated the advantages of the multi-task framework. We follow the similar framework in Fu et al. (2023); Hsieh et al. (2023); Magister et al. (2022), where the label prediction task and rationale task are applied to the training of the smaller language model. Specifically, we use different prompts to enforce the smaller language model to generate different types of output. Given an input question, for the label prediction task, the smaller language model outputs the prediction label while for the rationale task, it outputs the explanation about the prediction label.
3.2 Generation of contrastive knowledge
In what follows, we introduce the generation of the contrastive knowledge for dsitillation in detailed.
3.2.1 Extract diverse rationales from LLMs
We use Chain-of-Thought (CoT) prompting Wei et al. (2022b) to elicit and extract rationales from LLMs. As illustrated in Fig. 3, the LLM is provided with few examples to follow the output format. Instead of only generating one output for each input, we replace the “greedy decode” in CoT prompting with sampling from the language model’s decoder to generate a diverse set of reasoning paths Wang et al. (2022). We apply temperature sampling Renze and Guven (2024) and for each input, there are $K$ pairs of rationales and corresponding labels.
Figure 3: A case of the prompt and rationale output.
3.2.2 Build positive and negative rationale sets
Language models are not perfect reasoners and they might occasionally produce an incorrect reasoning path or make a mistake in one of the reasoning steps. In general, evidence shows that correct reasoning processes, even though they are diverse, tend to have greater agreement in their final answers than incorrect reasoning processes Wang et al. (2022). We thus choose the rationales with the most consistent label in all outputs of the LLM as positive rationale samples, and the remaining rationales are viewed as negative rationale samples. However, the sampling times mainly influence the proportion of positive and negative samples where the number of negative samples is always smaller), along with the time and storage costs.
To deal with this issue, we conduct a self-play mechanism that the student model generates its own rationales by sampling from its previous iterations during the training (details can be seen in Appendix A.1), and we regard these low-quality rationales as negative samples based on the hypothesise that the rationale quality of LLMs is higher than that of smaller models. As a result, for each input question $\mathbf{x}=[x_{1},x_{2},...,x_{n}]$, we collect a positive rationale set $S_{pos}=\{\mathbf{r}_{1}^{pos},\mathbf{r}_{2}^{pos},...,\mathbf{r}_{m}^{pos}\}$ and a negative rationale set $S_{neg}=\{\mathbf{r}_{1}^{neg},\mathbf{r}_{2}^{neg},...,\mathbf{r}_{k}^{neg}\}$.
3.3 Contrastive knowledge distillation
In this subsection, we represents our designed quality guided contrastive rationale distillation for better knowledge learning from the perspective of training process.
3.3.1 Train a discriminator to judge rationales
The quality of rationales for the same question still differs. Moreover, as the training epoch increases, the rationales generated by the above self-play may become gradually closer to the positive rationales, and then viewing them as negative samples is no longer reasonable. For this reason, there is a need to train a discriminator $\mathcal{D}$ that can effectively judge the positive and negative rationales and output a score that represents the quality of each rationale. The input of the discriminator $\mathcal{D}$ is the question and the rationale, and we take an encoder architecture to measure the score, i.e.,
$s_{j}^{pos}=\mathcal{D}\left(\mathbf{x},\mathbf{r}_{j}^{pos}\right)\;\;\mbox{or}\;\;s_{j}^{neg}=\mathcal{D}\left(\mathbf{x},\mathbf{r}_{j}^{neg}\right).$ (1)
We pretrain the $\mathcal{D}$ with the positive and negative rationales from the LLM, and during the training, the discriminator $\mathcal{D}$ is updated at regular epoch intervals (details can be seen in Appendix A.2), as the loss function can be formulated as
$\mathcal{L}_{\mathcal{D}}=\mathbb{E}_{\mathbf{x}}\left[-\log\frac{\sum_{j=1}^{m}\exp(s^{pos}_{j})}{\sum_{j=1}^{k}\exp(s^{neg}_{j})}\right].$ (2)
Figure 4: Illustration of the MinMax scheme for contrastive distillation.
3.3.2 Quality guided contrastive distillation
As mentioned in sec. 3.2.2, there is a diverse set of positive rationales. In addition, the negative rationales are of significance, which can enforce the smaller model away from their distribution. Since some of the negative samples are generated by the previous-iteration smaller model, the smaller model can further refine its reasoning capability through playing against instances of itself and promote the generated rationales closer to golden rationales of the LLM. Therefore, we propose a many-to-one contrastive distillation loss, while previous studies typically utilize a single rationale for each question and distill it into the smaller model, i.e.,
$\mathcal{L}_{cl}=\frac{1}{N}\sum_{i=1}^{N}[l(f(\mathbf{x}_{i}),S_{pos}^{i})-\beta\cdot l(f(\mathbf{x}_{i}),S_{neg}^{i})],$ (3)
where $\mathbf{x}_{i}$, $S_{pos}^{{\color[rgb]{0,0,0}i}}$, and $S_{neg}^{{\color[rgb]{0,0,0}i}}$ denote the $i$-th question, and the corresponding positive and negative rationale sets, respectively, $N$ is the number of questions, $\beta>0$ is a tunable hyper-parameter, the function $l(\cdot)$ denotes the cross entropy loss, and the function $f(\cdot)$ denotes the rationale generation for its given input. Here,
$l(f(\mathbf{x}_{i}),S_{pos}^{i})=\mathop{\min}_{\mathbf{r}_{j}^{pos,i}\in S_{pos}^{i}}\left\{l\left(f(\mathbf{x}_{i}),\mathbf{r}_{j}^{pos,i}\right)\right\},$ (4)
$l(f(\mathbf{x}_{i}),S_{neg}^{i})=\mathop{\max}_{\mathbf{r}_{j}^{neg,i}\in S_{neg}^{i}}\left\{l\left(f(\mathbf{x}_{i}),\mathbf{r}_{j}^{neg,i}\right)\right\},$ (5)
which are motivated by the sequence-level distillation Kim and Rush (2016) and illustrated in Fig. 4. Moreover, we set a margin $\delta$ for the negative rationales to balance the loss, i.e., $l(f(\mathbf{x}),\mathbf{r}_{j}^{neg})=\max(\delta-l(f(\mathbf{x}),\mathbf{r}_{j}^{neg}),0)$ with respect to the $j$-th negative rationale for an input question $\mathbf{x}$. Let us rethink the effectiveness of negative rationales generated by the previous-iteration smaller model, which enforces the smaller model to break out of local optima and yield a golden rationale that is closer to the output of the LLM. However, when the smaller model comes to converging, the previous-iteration smaller model is likely to output the rationales that are similar to those of the LLM, and then regarding them as negative samples is inaccurate. To address this issue, we introduce the quality guided distillation to optimize the training process and redefine the loss formulas in (4) and (5) as, respectively,
$l(f(\mathbf{x}_{i}),S_{pos}^{i})=s^{pos,i}\cdot\mathop{\min}_{\mathbf{r}_{j}^{pos,i}\in S_{pos}^{i}}\left\{l\left(f(\mathbf{x}_{i}),\mathbf{r}_{j}^{pos,i}\right)\right\},$ (6)
$l(f(\mathbf{x}_{i}),S_{neg}^{i})=(1-s^{neg,i})\cdot\mathop{\max}_{\mathbf{r}_{j}^{neg,i}\in S_{neg}^{i}}\left\{l\left(f(\mathbf{x}_{i}),\mathbf{r}_{j}^{neg,i}\right)\right\},$ (7)
where $s^{pos,i}$ and $s^{neg,i}$ are the corresponding quality scores obtained by the discriminator $\mathcal{D}$. It is reasonable that the positive rationales of higher quality should have larger weights across the datasets, while for the negative rationales of higher quality, it is on the contradiction. In the latter subsection 5.1, we will further discuss different schemes for the many-to-one distillation.
3.3.3 Training loss
The final training loss is given by
$\mathcal{L}_{total}=\alpha_{1}\mathcal{L}_{pred}+\alpha_{2}\mathcal{L}_{cl}+\alpha_{3}\mathcal{L}_{\mathcal{D}},$ (8)
where $\left\{\alpha_{i}\right\}_{i=1}^{3}>0$ are tunable hyper-parameters, $\mathcal{L}_{pred}$ represents the cross entropy loss of the label prediction task, $\mathcal{L}_{cl}$ is the many-to-one contrastive distillation loss in (3), and $\mathcal{L}_{\mathcal{D}}$ is the discriminator loss in (1).
4 Experiments
4.1 Experimental setting
Datasets. To validate QCRD, we conducted extensive experiments on four widely-used benchmark datasets (see details in Appendix Table 1) across three different natural language processing tasks, including SVAMP Patel et al. (2021) for arithmetic word problem solving, CQA Talmor et al. (2018) for commonsense question answering, as well as e-SNLI Camburu et al. (2018) and ANLI Nie et al. (2019) for natural language inference. The rationales we used were generated by GPT-3.5-turbo https://platform.openai.com/docs/models and open-sourced by Hsieh et al. (2023).
Implementation details. Following the properties of CoT and the comparative experimental studies in Hsieh et al. (2023), our QCRD adopted T5-base (220M parameters) and T5-small (60M parameters) as the smaller (student) model, respectively. $\alpha_{1}$ and $\alpha_{2}$ were set to $0.5$, and $\alpha_{3}$ was initially set to $0.5$ and multiplied by $0.9$ in each iteration. We also set $\beta=0.2$ and $\delta=4$. We sampled the LLM’s output five times with the temperature parameter being $0.7$, and sampled a five-iteration-before model with the temperature parameter being $1.5$. We trained our models on one A100-80G GPU and the batchsize is 32.
Baselines. We followed the comparison methods in Hsieh et al. (2023). Specifically, three methods in learning task-specific models were compared, i.e., (1) Finetuning, which is the standard finetuning with the prevailing pretrain-then-finetune paradigm that finetunes a model with ground-truth labels via standard label supervision Howard and Ruder (2018); 2) Single-Task, where the student model is distilled to predict labels with the teacher model’s predicted labels; 3) DSS Hsieh et al. (2023), where the smaller model is distilled with both the predict labels and rationales of the LLM.
Figure 5: Comparisons with varying sizes of training datasets on T5-base model.
4.2 Experimental results
Experiments across four benchmarks. We conducted expriments across four benchmarks with two types of T5-model to evaluate the effectiveness of our proposed method. In Table 1, we summarized the experimental results of the T5-base model distilled by our method and the baselines individually on all the four datasets. Of note, in Single-Task, the rationale and label were combined into a single sequence, which was then treated as the target during the training Hsieh et al. (2023). One can readily see from Table 1 that our method achieved superior performance over the baselines on all the four datasets.
Table 1: CoT distillation results on the T5-base model.
Model SVAMP CQA ANLI1 ESNLI
Finetuning 63.0 62.19 43.58 88.38
Single-Task 59.0 63.11 47.90 88.77
DSS 65.5 63.23 52.80 90.09
Ours 69.0 63.64 54.00 90.15
In like manner, we performed our method and the baselines individually on the T5-small model, and their performance on all the four datasets was presented in Table 2. Our method consistently surpassed the baselines on all the four datasets as expected.
Table 2: CoT distillation results on the T5-small model.
Model SVAMP CQA ANLI1 ESNLI
Finetuning 45.00 43.16 42.00 82.90
DSS 48.00 45.21 42.80 84.23
Ours 50.50 46.11 44.10 85.30
Distillation with LLM labels. To evaluate the impact of label qualities on CoT distillation, without loss of generality, we conducted additional experiments on the three datasets (namely, CQA, ANLI1, and ESNLI) using the T5-base model distilled by our method and DSS. Instead of using ground truth labels, we employed the labels generated by GPT-3.5-Turbo to distill the student models. The results were presented in Table 3. On one hand, from the top of Table 3, it demonstrates the effectiveness of temperature sampling and self-consistency, which help denoise rationales and their corresponding labels. On the other hand, the results at the bottom of Table 3 indicate that our method outperforms DSS on CQA and ANLI1, even when utilizing the labels generated by the noise-affected LLM. It suggests that our method is more robust and can yield improved performance.
Table 3: CoT distillation results on the T5-base model.
ModelP CQA ANLI1 ESNLI
GPT 3.5 66.30 78.21 66.27
GPT 3.5 with SC 69.05 80.15 67.08
DSS 59.15 44.10 74.88
Ours 59.80 46.70 74.88
Distillation with smaller datasets. In addition, to demonstrate the superiority of our method on smaller datasets, we compared the performance of DSS and our method using T5-base models on varying sizes of each of the four datasets. Fig. 5 illustrated that our method achieved better performance in most cases. It indicate the robustness and generality of our proposed QCRD.
Ablation study on our QCRD. Compared to previous related methods, the contrastive distillation in our QCRD introduces several key enhancements as follow. (1) The extension and denoising for positive knowledge (ED): we sample the output of the LLM and leverage the self-consistency to denoise rationales. (2) The distillation for negative knowledge (NK): we incorporate a self-play mechanism to generate low-quality rationales as negative samples. (3) The guidance of the Quality Judge (QJ): the use of a quality judgement discriminator helps assess rationales and optimize the training process. Additional experiments were so conducted on SVAMP to evaluate the effectiveness of each module, with the results being summarized in Table 4. The findings demonstrate that integrating more high-quality rationales significantly improves performance, while the inclusion of negative rationales proves effective. The discriminator mechanism plays a positive role by considering the quality of each rationale, and we further find that the results when using the Quality Judge are more stable.
Table 4: Ablation study on T5-base model.
w/o ED w/o NK w/o QJ SVAMP
65.5
67.0
68.5
69.0
5 Discussion
5.1 Different schemes for contrastive distillation
In sec. 3.3.2, we define the many-to-one distillation by taking the min loss for the positive rationales and max loss (after margin operation) for the negative rationales, which imposes a relatively weak constraint on rationale alignment. We here further discuss different schemes for the many-to-one distillation. (1) MaxMin: we compute the max loss for the positive rationales and min loss for the negative rationales. This scheme enforces the smaller model to learn hard rationale examples. (2) Sampling: we randomly choose a positive rationale and a negative rationale for each anchor. (3) Mean: we average the loss for all rationales. (4) Weighted mean (W-mean): we weight the loss with quality scores and then average the loss. The results of the T5-base model distilled by our method on SVAMP were presented in Table 5 with respect to the above different schemes. One can clearly see that MinMax achieved the best performance. Besides, the scheme of Mean had a negative impact on the results. The reason may be that the positive and negative rationales generated by temperature sampling may be of great difference, and especially for the positive rationales, enforcing the small model align with multi-target outputs is not suitable.
Table 5: Results of our method with different many-to-one distillation schemes on SVAMP.
Model MinMax MaxMin Sampling Mean W-mean
T5-base 69.0 67.0 66.5 65.0 66.0
5.2 Influence of the sampling count
In the above experiments, we sampled the output of the LLM five times and the output of iteration-before model once. We further explore the influence of the sampling count. When fixing the sampling counts for the iteration-before model, the results of setting different sampling counts for the LLM on SVAMP were displayed in the top of Table 6. Moreover, when fixing the sampling counts for the LLM, the results on SVAMP were displayed in the bottom of Table 6 in terms of different numbers of generated negative samples. We found that sampling many negative rationales had an adverse impact on the performance, and the best performance was achieved when $k$ was 5. Note that when $k=1$, the performance of our method was still better than that of other related methods, again indicating the effectiveness of the use of negative rationales.
Table 6: Results of our method on SVAMP with different sampling counts.
Positive sample m 1 5 10 20
T5-base 67.5 69.0 68.0 68.5
negative sample k 0 1 2 3
T5-base 67.0 69.0 68.5 66.0
5.3 Rational analysis of negative rationales by self-play
Temperature sampling is a commonly used decoding strategy for LLMs’ generation process. In our method, we generated low-quality rationales by sampling the iteration-before smaller models. As displayed in Appendix Figure 2, we provided a case of the output rationale with different temperature settings for better illustration. The golden rationale can be obtained by greedy search; however, when the temperature was set to a higher value, the output rationales are likely various but inaccurate.
The results of the T5-base model distilled by our method on SVAMP were displayed in Table 7 in terms of different negative sampling temperature settings. It was observed that when no sampling was performed (i.e., $\tau=0$) or a lower temperature value was used (i.e., $\tau=0.7$), the smaller model exhibited relatively poorer performance and showed larger fluctuations in accuracy. The best results were achieved when the temperature $\tau$ was set to $1.5$. The reason for this can be attributed to the fact that when the model approaches convergence, the output rationales with lower temperature values tend to be similar to the golden ones. Considering these similar outputs as negative samples can lead to detrimental effects. All in all, we sampled the output rationales of iteration-before models several times and utilized these generated low-quality rationales as negative samples to enforce the small models away from these distributions and boost the performance.
Table 7: Results of our method on SVAMP with different negative sampling temperature settings.
$\tau$ 0 0.7 1 1.5 2
T5-base 65.0 64.5 67.5 69.0 67.5
5.4 The quality assessment for the output rationales
We assessed the quality of CoT examples using GPT-3.5-turbo. Inspired by the ranking model, we prompted GPT-3.5-turbo to rank the rationales generated by both DSS and our QCRD, rather than providing scores based on the quality of the rationales. This is easier for the LLM and allows for a more straightforward comparison. The prompt provided to GPT-3.5-turbo was presented in Table 8. To evaluate the rationales, we randomly selected $50$ examples from each of the four datasets and asked GPT-3.5-turbo to determine rationales of which our method was better. We then aggregated the counts of "DSS is better," "Both are good," and "QCRD is better," as shown in Table 9. From the results, we observed that on SVAMP, CQA, and ESNLI, the model trained using our method generated better rationales than using DSS. However, on ANLI1, the model trained using DSS exhibited slightly better performance.
Table 8: The prompt for GPT-3.5-turbo to judge rationales.
The prompt for GPT-3.5-turbo
There is an input pair of a question and an answer of a taskname task, and we provide you two explanations. You need judge which explanation is better. The better explanation should be more accurate and explain the answer better.
Table 9: The quality assessment results on T5-base model for different sampling temperature settings, where three numbers represent the counts of "DSS is better", "Both are good", and "QCRD is better", respectively.
SVAMP CQA ANLI1 ESNLI
$\tau=0$ 21/0/29 19/6/25 26/1/23 17/11/22
$\tau=0.7$ 22/0/28 20/6/24 25/1/24 14/3/33
6 Conclusion
The knowledge distillation of CoT rationales from LLMs into smaller reasoners using a multi-task framework has been empirically shown to enhance performance. Building upon this framework, this paper introduces general CoT distillation method, incorporating a contrastive learning perspective that considers both positive and negative knowledge. To generate positive and negative rationales, we propose an innovative approach that combines temperature sampling, the self-consistency of LLMs, and the self-play of the small reasoners themselves. Additionally, we develop a many-to-one contrastive distillation loss for better knowledge learning, where an online-update discriminator is used to judge the qualities of rationales and assign weights for optimizing the training process across the whole datasets. Extensive experiments conducted on multiple reasoning tasks demonstrate the superiority of our method over the previous ones.
Appendix A Appendix
Table 1: Descriptions of the four benchmark datasets.
Dataset Training Validation Test
SVAMP 720 80 200
CQA 8766 975 1221
ANLI1 16946 1000 1000
ESNLI 549367 9842 9824
A.1 Iteration-before-models for negative rationale generators
In this paper, we dynamically generated negative rationales using iteration-before-models through online temperature sampling. As depicted in Fig. 1, to select a j-iteration-before-model for the negative rationale generator, we need to save a minimum of $j$ checkpoints for the model. This allows us to train the model end-to-end instead of using a multi-turn approach.
Figure 1: A case of the j-iteration-before-model for negative rationale generator.
A.2 Details about the Quality Judge
We incorporated a discriminator into our training process to assess the quality of rationales and assign corresponding weights to the losses. To construct the discriminator, we leveraged the encoder of the T5-base model along with two additional linear layers to compute the quality score. Prior to training, the discriminator needs to be pretrained using the output rationales generated by LLMs with applying data augmentations to the negative rationales. Specifically, we employed word replacement with the assistance of StanfordNLP Zeman et al. (2018). Additionally, throughout the training process, we dynamically updated the discriminator using a decayed learning rate for better learning the negative cases from iteration-before-models.
Figure 2: A case study of the output rationale on SVAMP with different temperature sets.